A parallel-plate capacitor has plates with an area of 250 cm2 separated by 1.0 mm. What is the electric field between the plates when they are charged with 7.000×10-7 C?
Answer #1

Solution:

Using the formula,

=> E = Q/(e0 * A)

=> E = 7 * 10^-7 /(8.85 * 10^-12 * 250 * 10^-4)

=> E = 7 * 10^9 /(8.85 * 250)

=> E = 3.16 * 10^6 N/C

Please comment for queries please rate

Thanks